# Barking Up The Syntactic Tree: Enhancing VLM Training with Syntactic Losses

### v1

@@ 3 Method
Algorithm 1 Data Pair Creation: Noun Phrases Generation and Subject Constraints
@@
3:     $\{(\mathbf{C}_{1}^{k},S(\mathbf{C}_{1}^{k})),\dots,(\mathbf{C}_{N}^{k},S(\mathbf{C}_{N}^{k}))\}\leftarrow$ LLM$(\text{cap}_{k})$triangleright Generate noun phrases and their subject
4:     for each phrase $\mathbf{C}_{i}^{k}$ in $\{\mathbf{C}_{1}^{k},\dots,\mathbf{C}_{N}^{k}\}$ do
5:         $\text{Candidates}\leftarrow\{\mathbf{C}_{j}^{k}\mid j\neq i\land S(\mathbf{C}_{j}^{k})\neq S(\mathbf{C}_{i}^{k})\}$
6:         if $\text{Candidates}\neq\emptyset$ then
7:              Randomly select $\mathbf{C}_{j}^{k}$ from Candidates
8:              $\mathcal{P}\leftarrow\mathcal{P}\cup\{(\mathbf{I}_{k},\mathbf{C}_{i}^{k},\mathbf{C}_{j}^{k})\}$
9:         end if
10:     end for
11:end for
12:return $\mathcal{P}$
3.1 Base Model and Objectives
Our formulation, HIST, is applicable to any VLM model that is capable of obtaining spatial image-text attention. We illustrate its efficacy on two popular VLM models: ALBEF [24] and BLIP [25]. Both of these models consist of an unimodal image encoder $\boldsymbol{\phi_{v}}$, a unimodal text encoder $\boldsymbol{\phi_{t}}$, and a multimodal encoder $\boldsymbol{\phi_{m}}$. The output of the image encoder and text encoder are represented as $\mathbf{V}=\{\mathbf{v}_{cls},\mathbf{v}_{1},\cdots,\mathbf{v}_{N_{v}}\}\in\mathbb{R}^{(N_{v}+1)\times d_{v}}$, where $N_{v}$ is the number of visual tokens and $d_{v}$ is the visual embedding dimension, and $\mathbf{T}=\{\mathbf{t}_{cls},\mathbf{t}_{1},\cdots,\mathbf{t}_{N_{t}}\}\in\mathbb{R}^{(N_{t}+1)\times d_{t}}$ where $N_{t}$ is the number of text tokens and $d_{t}$ is the text embedding dimension. Both models utilize the three commonly used objectives for learning visual and textual representations: image-text contrastive (ITC), image-text matching (ITM), masked language modeling (MLM). We provide a brief overview of these objectives here as we also leverage them during our training.
Image-Text Contrastive Learning (ITC). The ITC objective aims to align unimodal representations by maximizing similarity for matching image-text pairs and minimizing it for non-matching pairs. For an image $\mathbf{I}$ and text $\mathbf{T}$, we define the similarity as:
$s(\mathbf{I},\mathbf{T})=g_{v}(\mathbf{v}_{\text{cls}})^{\top}g_{t}(\mathbf{t}_{\text{cls}}),$
where $g_{v}$ and $g_{t}$ are linear projections of the [CLS] token from the image patches embedding and the text embedding.
The image-to-text and text-to-image similarities are calculated as:
$p^{i2t}_{m}(\mathbf{I})=\frac{\exp(s(\mathbf{I},\mathbf{T}_{m})/\tau)}{\sum_{m=1}^{M}\exp(s(\mathbf{I},\mathbf{T}_{m})/\tau)},$
$p^{t2i}_{m}(\mathbf{T})=\frac{\exp(s(\mathbf{T},\mathbf{I}_{m})/\tau)}{\sum_{m=1}^{M}\exp(s(\mathbf{T},\mathbf{I}_{m})/\tau)},$
where $\tau$ is a temperature parameter and M being the image or text queue following MoCo [13]. The ITC loss is defined as follows:
$\mathcal{L}_{\text{itc}}=\frac{1}{2}\mathbb{E}_{(\mathbf{I},\mathbf{T})\sim\mathcal{D}}\Big{[}\mathcal{H}\big{(}y^{i2t}(\mathbf{I}),p^{i2t}(\mathbf{I})\big{)}+$
$\hskip 137.31255pt+\mathcal{H}\big{(}y^{t2i}(\mathbf{T}),p^{t2i}(\mathbf{T})\big{)}\Big{]},$
where $y^{i2t}(I)$ and $y^{t2i}(T)$ are one-hot ground-truth vectors, $\mathcal{H}$ denotes the cross-entropy, and $\mathbb{E}$ is expectation over the training dataset $\mathcal{D}$ in image-text pairs.
Image-Text Matching (ITM). The ITM objective predicts whether an image-text pair $(\mathbf{I},\mathbf{T})$ is positive (matched) or negative (not matched). This is achieved by passing the fused [CLS] embedding from the multimodal encoder through a fully connected layer followed by a softmax to predict a binary probability $p^{\text{itm}}(\mathbf{I},\mathbf{T})$. The ITM loss $L_{\text{itm}}$ is defined as:
$\mathcal{L}_{\text{itm}}=\mathbb{E}_{(\mathbf{I},\mathbf{T})\sim\mathcal{D}}\,\mathcal{H}\left(y^{\text{itm}},p^{\text{itm}}(\mathbf{I},\mathbf{T})\right),$
where $y^{\text{itm}}$ is a two-dimensional one-hot vector representing the ground-truth label (matched or not).
Masked Language Modeling (MLM). Masked Language Modeling (MLM) integrates contextual information from both the text and the paired image to predict masked tokens within the text. Tokens in the input text are randomly masked out and replaced with a special [MASK] token. Let $\hat{\mathbf{T}}$ denote the masked text and $p^{\text{msk}}(\mathbf{I},\hat{\mathbf{T}})$ denote the model’s predicted probability distribution for the masked token. The MLM objective $\mathcal{L}_{\text{mlm}}$ minimizes the cross-entropy loss:
$\mathcal{L}_{\text{mlm}}=\mathbb{E}_{(\mathbf{I},\hat{\mathbf{T}})\sim\mathcal{D}}\,\mathcal{H}\left(y^{\text{msk}},p^{\text{msk}}(\mathbf{I},\hat{\mathbf{T}})\right),$
where $y^{\text{msk}}$ is a one-hot distribution representing the ground-truth token for the masked position.
Certain VLMs though, like BLIP [25], use a language modeling loss focused on predicting the next word token rather than randomly masking words within sentences and predicting the masked tokens. Essentially, this approach uses a cross-entropy loss for next-word prediction.
@@ 3.2 HIerarchically STructured (HIST) Losses
To enhance cross-modal alignment at different levels of granularity, we introduce two additional sets of hierarchically structured losses which we describe in this section; mainly the Subject and the Addition loss.
@@
Phrase parsing. As illustrated in Algorithm 1, we follow SelfEQ [14] and perform phrase chunking using Vicuna [7]. Specifically, in-context examples are provided to Vicuna [7] to first generate noun phrases from caption, and then each noun phrase is re-inputted to Vicuna to detect the subject. To construct composite phrase dataset $\mathcal{P}$, for each noun phrase $\mathbf{C}_{i}^{k}$, we identify potential pairing candidates $\{\mathbf{C}_{j}^{k}\}$ from other noun phrases from the same image but not referring to the same subject (these do not necessarily come from the same caption). We then randomly select a pairing candidate and form a data pair $(\mathbf{I}_{k},\mathbf{C}_{i}^{k},\mathbf{C}_{j}^{k})$, and add it to $\mathcal{P}$. This process repeats for all phrases in each caption, ultimately generating $\mathcal{P}$, which contains the selected image-composite caption phrase pairs for further use.
@@
Phrase Loss. With the noun phrases we obtained above, we apply the ITC, ITM and MLM on the phrases directly: $\mathcal{L}_{phrase}=$
$\mathcal{L}_{ITC}(\{(\mathbf{I},{\mathbf{C}_{i}})\})+\mathcal{L}_{ITM}(\{(\mathbf{I},{\mathbf{C}_{i}})\})+\mathcal{L}_{MLM}(\{(\mathbf{C}_{i})\}).$
@@
Subject Loss. The motivation of the Subject loss comes from the fact that a potentially much longer noun phrase, at the phrase level, may obfuscate text-image alignment by focusing more on adjectives (that maybe common among many phrases) as compared to the noun that tends to be more semantically important. With such an observation, we add a Subject Loss during training, which enforces consistency, aligning the subject of the phrase with the corresponding visual representation. Specifically, for a phrase $\mathbf{C}_{i}$ with main subject $\mathbf{S}_{i}$, the loss is:
$\mathcal{L}_{subject}=\mathcal{L}_{ITC}(\{(\mathbf{I},{\mathbf{S}_{i}})\})+\mathcal{L}_{ITM}(\{(\mathbf{I},{\mathbf{S}_{i}})\}).$
where $\mathcal{L}_{ITC}$ and $\mathcal{L}_{ITM}$ are defined above.
@@
Localization in VLMs. We first introduce how we obtain localization map for phrases and then discuss the novel Addtion loss we proposed.
@@
The text prompt and the image first go through the modality-specific encoders $\boldsymbol{\phi_{t}}$ and $\boldsymbol{\phi_{v}}$ respectively, then the respective embeddings are input into the cross-attention fusion module. For each cross-attention layer, the text encodings serve as query vectors and the image patch encodings are as key and value vectors. From a particular cross-attention layer (and specific head), we extract the corresponding text-to-image cross-attention map as shown below. Note that different layers and heads lead to drastic performance differences, and we follow [14, 16] for the best layer-head combination; resulting in
@@
$\mathbf{A}_{\mathbf{C}_{i}}=\texttt{Softmax}\left(\frac{\boldsymbol{\phi_{t}}(\mathbf{T})\ \mathbf{W}_{Q}(\boldsymbol{\phi_{v}}(\mathbf{I})\mathbf{W}_{K})^{\top}}{\sqrt{D_{t}}}\right),$
@@
where $\mathbf{A}_{\mathbf{C}_{i}}$ has shape $(K+1)\times P\times P$ with $P$ being the number of image patches and $K$ being the number of text tokens. To get the phrase specific attention map, we take the mean over all K tokens and get final map of size $P\times P$. This map can serve as proxy for segmentation [16] and can further be enhanced by GradCAM [34], to highlight discriminative regions, as shown in [39, 16]. Applying this technique results in the final localization map:
@@
$\mathbf{G}_{\mathbf{C}_{i}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{i}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},i}}{\partial\mathbf{A}_{\mathbf{C}_{i}}}\right),$
@@
Addition Loss. We are now ready to define our novel Addition loss at the combined phrase level. For a pair of phrases $\mathbf{C}_{i}$, $\mathbf{C}_{j}$, extracted from captions of the same image, and the combined composite phrase $\mathbf{C}_{ij}$ we extract the cross-attention maps and then the corresponding localization maps as follows:
$\mathbf{G}_{\mathbf{C}_{i}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{i}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},i}}{\partial\mathbf{A}_{\mathbf{C}_{i}}}\right),$
$\mathbf{G}_{\mathbf{C}_{j}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{j}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},j}}{\partial\mathbf{A}_{\mathbf{C}_{j}}}\right),$
$\mathbf{G}_{\mathbf{C}_{ij}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{ij}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},ij}}{\partial\mathbf{A}_{\mathbf{C}_{ij}}}\right).$
We observed that vision-language models often focus disproportionately on the most prominent objects within an image, which can lead to poor alignment. Therefore, we introduce the Addition loss to improve the model’s capacity to attend to multiple objects simultaneously. It is defined as:
$\mathcal{L}_{\text{add}}=\sum_{P}\sum_{P}\left|\mathbf{G}_{\mathbf{C}_{i}}+\mathbf{G}_{\mathbf{C}_{j}}-\mathbf{G}_{\mathbf{C}_{ij}}\right|.$
The loss enforces that the attention map for the combined phrase approximates the sum of individual attention maps, ensuring that the model considers all relevant objects.
3.3 Total Hierarchical Alignment Loss
The total loss combines all the losses mentioned above:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{phrase}}+\mathcal{L}_{\text{subject}}+\mathcal{L}_{\text{add}}.$
This hierarchical loss formulation allows our model to align fine-grained visual and textual semantics effectively, while maintaining global coherence across the image-text pair. Each component contributes to a specific level of alignment, from individual phrases to subjects and combined phrases, ultimately refining the model’s ability to handle complex, multi-object scenes and improving its localization ability.
@@ 4 Experiment
Implementation details. We apply HIST on ALBEF [24] and BLIP [25] and train the model with Visual Genome [21] and MSCOCO [28]. ALBEF combines a ViT-B model for encoding images and a BERTbase model for encoding text. Following SelfEQ [14], we train ALBEF with image size of $256\times 256$ with the compostie objective function $\mathcal{L}$ = $\alpha$ $\mathcal{L}_{\text{vl}}$ + (1 - $\alpha$) ($\mathcal{L}_{\text{vl}}$ + $\mathcal{L}_{\text{HIST}}$), where $\alpha$ is initially set to 0 and increments to 1, remaining constant after the second epoch. We train our model with with 4 NVIDIA A6000 with learning rate of 2e-5. For the addtion loss, we employ the head average gradcam map from the 3rd cross-attention layer. For BLIP, we train BLIPlarge, which adopts VIT-L/16 as the image encoder, BERT as the text encoder, and insert an extra cross-attention layer for each transformer block of BERT. Following BLIP [25], we train BLIP with image size of $224\times 224$, on 4 NVIDIA A6000 with learning rate 1e-5. For the addtion loss, we leverage the head average gradcam map from the 9th cross-attention layer. We will release all models, data and code upon acceptance of the paper.
@@ 4.1 Visual Grounding
Method Training Flickr30K Referit
InfoGround [12] VG-boxes 76.7 -
VMRM [9] VG-boxes 81.1 -
AMC [45] VG-boxes 86.6 73.2
ALBEF [24] Zero-shot 79.4 59.7
BLIP [25] Zero-shot 80.1 51.6
[36] VG 75.6 66.0
g++ [35] VG 80.0 70.3
ALBEF + SelfEQ [14] VG 81.9 67.4
(Ours) ALBEF + HIST VG 83.6 69.5
(Ours) BLIP + HIST VG 81.5 58.1
[36] COCO 75.4 61.0
g++ [35] COCO 78.1 61.5
ALBEF + SelfEQ [14] COCO 84.1 62.8
(Ours) ALBEF + HIST COCO 85.3 63.4
(Ours) BLIP + HIST COCO 84.7 57.6
@@
InfoGround [12] VG-boxes 39.8 44.1
VMRM [9] VG-boxes 58.9 50.3
AMC [45] VG-boxes 80.3 64.6
ALBEF [24] Zero-shot 69.4 53.8
BLIP [25] Zero-shot 70.2 52.4
ALBEF + SelfEQ [14] VG 74.1 55.1
ALBEF + SelfEQ [14] VG 75.1 55.5
(Ours) ALBEF + HIST VG 74.7 56.7
(Ours) BLIP + HIST VG 75.3 56.4
@@
Figure 3: Qualitative Result for ALBEF, SelfEQ and HIST + ALBEF on Visual Grounding and Referring Segmentation. Images are from RefCOCO+. The enlarged red stars represent the top 4 locations with the highest predicted attention value from respective methods. To obtain segmentation mask, we input the four points into SAM [19] as point prompts. We note that the HIST accurately detect all objects but the mask can over-segment or under-segment.
Method Training CIoU
ALBEF [24] Zero-shot 29.4
ALBEF + SelfEQ [14] COCO 30.3
(Ours) ALBEF + HIST COCO 35.6
Table 3: Complex Referring Segmentation on RefCOCO+ [46].

### v2

@@ 3.2 Base Models and Objectives
Algorithm 1 Data Pair Creation: Phrases Generation and Subject Constraints
@@
3:    $\{(\mathbf{C}_{1}^{k},S(\mathbf{C}_{1}^{k})),\dots,(\mathbf{C}_{N}^{k},S(\mathbf{C}_{N}^{k}))\}\leftarrow$ LLM$(\text{cap}_{k})$
4$\triangleright$ Generate phrases and their subject
5:    for each phrase $\mathbf{C}_{i}^{k}$ in $\{\mathbf{C}_{1}^{k},\dots,\mathbf{C}_{N}^{k}\}$ do
6:         $\text{Candidates}\leftarrow\{\mathbf{C}_{j}^{k}\mid j\neq i\land S(\mathbf{C}_{j}^{k})\neq S(\mathbf{C}_{i}^{k})\}$
7:         if $\text{Candidates}\neq\emptyset$ then
8:             Randomly select $\mathbf{C}_{j}^{k}$ from Candidates
9:             $\mathbf{C}_{ij}^{k}\leftarrow\text{Concatenate}(C_{i}^{k},and,C_{j}^{k})$
10:             $\mathcal{P}\leftarrow\mathcal{P}\cup\{(\mathbf{I}_{k},\mathbf{C}_{i}^{k},\mathbf{C}_{j}^{k},\mathbf{C}_{ij}^{k})\}$
11:         end if
12:    end for
13:end for
14:return $\mathcal{P}$
@@ 3.3 HIerarchically STructured (HIST) Losses
We first parse image captions into hierarchical structures with three levels using an LLM. After that, we propose different losses on each level and across levels.
@@
Phrase parsing. Following SelfEQ [16] we identify phrases in image captions using an in-context-learning Vicuna [8]. Specifically, we first identify phrases from each caption, and one subject from each phrase. To create composite phrase dataset $\mathcal{P}$, for each phrase $\mathbf{C}_{i}^{k}$, we identify potential pairing candidates $\{\mathbf{C}_{j}^{k}\}$ from other phrases from the same image but do not have the same subject identified by Vicuna (these do not necessarily come from the same caption). We then randomly select a pairing candidate $\mathbf{C}_{j}^{k}$ for $\mathbf{C}_{i}^{k}$, and form a composite phrase $\mathbf{C}_{ij}^{k}$ by connecting two phrases using the conjunction “and”. The data tuple $(\mathbf{I}_{k},\mathbf{C}_{i}^{k},\mathbf{C}_{j}^{k},\mathbf{C}_{ij}^{k})$ is then added to $\mathcal{P}$. This process is repeated for all phrases in each caption, ultimately generating $\mathcal{P}$, which contains the selected image, phrase pairs, and a composite phrase for further use.
@@
Phrase Loss. With the phrases we obtained above, we apply the ITC, ITM and MLM on the phrases directly: $\mathcal{L}_{phrase}$
$=\mathcal{L}_{\text{itc}}\left(\{(\mathbf{I},\mathbf{C}_{i})\}\right)+\mathcal{L}_{\text{itm}}\left(\{(\mathbf{I},\mathbf{C}_{i})\}\right)+\mathcal{L}_{\text{mlm}}\left(\{\mathbf{C}_{i}\}\right)$ (1)
@@
Subject Loss. The motivation of the Subject loss comes from the fact that a potentially much longer phrase, at the phrase level, may obfuscate text-image alignment by focusing more on adjectives (that maybe common among many phrases) as compared to the object or actor that tends to be more semantically important. Relying on this intuition, we add a Subject Loss during training, which enforces consistency, aligning the subject of the phrase with the corresponding visual representation. Specifically, for a phrase $\mathbf{C}_{i}$ with main subject $\mathbf{S}_{i}$, the loss is:
$\mathcal{L}_{\text{subject}}=\mathcal{L}_{\text{itc}}\left(\{(\mathbf{I},\mathbf{S}_{i})\}\right)+\mathcal{L}_{\text{itm}}\left(\{(\mathbf{I},\mathbf{S}_{i})\}\right)$ (2)
where $\mathcal{L}_{itc}$ and $\mathcal{L}_{itm}$ are defined above. Note that we follow the original $\mathcal{L}_{itc}$ and incorporate MoCo [15] to account for the potential positives in the negative pairs.
@@
Localization in VLMs. We first introduce how we obtain localization map for phrases and then discuss the novel Composition loss we propose.
@@
The text prompt and image are first processed by respective modality encoders, $\boldsymbol{\phi_{t}}$ and $\boldsymbol{\phi_{v}}$. The resulting embeddings are then passed into the cross-attention fusion module. For each cross-attention layer, the text embeddings serve as query vectors and the image patch embeddings are as key and value vectors. From a particular cross-attention layer (and specific head), we extract the corresponding text-to-image cross-attention map as shown below. Note that different layers and heads lead to drastic performance differences, and we follow [16, 18] for the best layer-head combination; resulting in
@@
$\mathbf{A}_{\mathbf{C}_{i}}=\texttt{Softmax}\left(\frac{\boldsymbol{\phi_{t}}(\mathbf{T})\ \mathbf{W}_{Q}(\boldsymbol{\phi_{v}}(\mathbf{I})\mathbf{W}_{K})^{\top}}{\sqrt{D_{t}}}\right)$ (3)
@@
where $\mathbf{A}_{\mathbf{C}_{i}}$ has shape $(K+1)\times P\times P$ with $P\times P$ being the number of image patches and $K$ being the number of text tokens. To get the phrase specific attention map, we take the mean over all K tokens and get final map of size $P\times P$. This map can serve as proxy for segmentation [18] and can further be enhanced by GradCAM [40], to highlight discriminative regions, as shown in [49, 18]. Applying this technique results in the final localization map:
@@
$\mathbf{G}_{\mathbf{C}_{i}}=\texttt{ReLU}\left(\mathbf{A}_{\mathbf{C}_{i}}\cdot\frac{\partial\mathcal{L}_{\text{ITM},i}}{\partial\mathbf{A}_{\mathbf{C}_{i}}}\right).$ (4)
@@
Composition Loss. We are now ready to define our novel Composition loss at the composite phrase level. For a pair of phrases $\mathbf{C}_{i}$, $\mathbf{C}_{j}$, extracted from captions of the same image, and the composite phrase $\mathbf{C}_{ij}$, we extract the cross-attention maps and then the corresponding localization maps $\mathbf{G}_{\mathbf{C}_{i}}$, $\mathbf{G}_{\mathbf{C}_{j}}$ and $\mathbf{G}_{\mathbf{C}_{ij}}$ as in equation 3 and 4.
We observed that vision-language models often focus disproportionately on the most prominent objects within an image, which can lead to poor alignment. Therefore, we introduce the Composition loss to improve the model’s capacity to attend to multiple objects simultaneously. It is defined as:
$\mathcal{L}_{\text{comp}}=\sum_{P}\sum_{P}\left|\mathbf{G}_{\mathbf{C}_{i}}+\mathbf{G}_{\mathbf{C}_{j}}-\mathbf{G}_{\mathbf{C}_{ij}}\right|.$ (5)
The loss enforces that the attention map for the composite phrase approximates the sum of individual attention maps, ensuring that the model considers all relevant objects.
Total Hierarchical Alignment Loss. The total loss combines all the losses mentioned above:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{phrase}}+\mathcal{L}_{\text{subject}}+\mathcal{L}_{\text{comp}}.$
This hierarchical loss formulation allows our model to align fine-grained visual and textual semantics effectively, while maintaining global coherence across the image-text pair. Each component contributes to a specific level of alignment, from individual phrases to subjects and composite phrases, ultimately refining the model’s ability to handle complex, multi-object scenes and improving its localization ability.
@@ 4 Experiment
Implementation details. We apply HIST on ALBEF [27] and BLIP [28] and train the model with Visual Genome [24] and MSCOCO [32]. ALBEF combines a ViT-B model for encoding images and a BERTbase model for encoding text. Following SelfEQ [16], we train ALBEF with image size of $256\times 256$ with the compostie objective function $\mathcal{L}$ = $\alpha$ $\mathcal{L}_{\text{vl}}$ + (1 - $\alpha$) ($\mathcal{L}_{\text{vl}}$ + $\mathcal{L}_{\text{HIST}}$), where $\alpha$ is initially set to 0 and increments to 1, remaining constant after the second epoch. We train our model with with 4 NVIDIA A6000 with learning rate of 2e-5. For the Composition loss, we employ the head average gradcam map from the 3rd cross-attention layer. For BLIP, we train BLIPlarge, which adopts VIT-L/16 as the image encoder, BERT as the text encoder, and insert an extra cross-attention layer for each transformer block of BERT. Following BLIP [28], we train BLIP with image size of $224\times 224$, on 4 NVIDIA A6000 with learning rate 1e-5. For the Composition loss, we leverage the head average gradcam map from the 9th cross-attention layer.
@@ 4.1 Visual Grounding
Method Training Flickr30K Referit AVG
[…]
        + HIST(Ours) VG 81.5 58.1 69.8
[…]
        + HIST(Ours) VG 83.6 69.5 76.6
[…]
        + HIST(Ours) COCO 84.7 57.6 71.2
[…]
        + HIST(Ours) COCO 85.3 63.4 74.4
@@
[…]
        + HIST(Ours) VG 75.3 56.4
[…]
        + HIST(Ours) VG 74.7 56.7
@@
Figure 3: Qualitative Result for ALBEF, /w SelfEQ and /w HIST on Visual Grounding and Referring Segmentation. Images are from RefCOCO+. The enlarged red stars represent the top 4 locations with the highest predicted attention value from respective methods. To obtain segmentation mask, we input the four points into SAM [22] as point prompts. We note that the HIST accurately detects all objects.